# FMCW Radar Principles and Human Activity Recognition Systems: Foundations, Techniques, and Applications

### v1

@@ 23.2 Conclusion
Generating heatmaps from Range-Doppler Maps and using CNNs for target detection is an effective approach for radar data processing [11]. Heatmaps provide a visual representation of radar data, and CNNs automatically extract spatial features from the heatmaps, enabling the classification of targets in real-time radar applications. With the ability to handle noise and extract meaningful features, CNNs play a critical role in enhancing radar-based object detection systems.
@@ 23.3.1 Using CNN for Heatmap Classification
17class SimpleCNN(nn.Module):
@@
20 self.conv1 = nn.Conv2d(1, 32, kernel_size=3, padding=1) # 1 input channel (grayscale), 32 output filters
21 self.pool = nn.MaxPool2d(2, 2)
22 self.conv2 = nn.Conv2d(32, 64, kernel_size=3, padding=1)
23 self.conv3 = nn.Conv2d(64, 64, kernel_size=3, padding=1)
24 self.fc1 = nn.Linear(64 * 25 * 25, 64) # Flatten after pooling
25 self.fc2 = nn.Linear(64, 2) # Binary classification (target or noise)
@@
42criterion = nn.CrossEntropyLoss() # For binary classification with logits
@@ 23.3.2 PointNet and PointNet++ for Point Cloud Processing
PointNet is a deep learning model that directly processes point cloud data without converting it into voxel grids or meshes [6]. The architecture of PointNet is designed to handle unordered sets of points, making it ideal for processing radar-generated point clouds. The model extracts global features from the entire point cloud by applying a series of shared multi-layer perceptrons (MLPs) to each point, followed by a max pooling operation to aggregate the features.
@@
PointNet++ builds upon the success of PointNet by introducing a hierarchical structure that captures both global and local features [20]. While PointNet extracts global features from the entire point cloud, PointNet++ divides the point cloud into smaller regions, applying PointNet to each region to extract local features. These local features are then combined to form a hierarchical representation of the point cloud, making PointNet++ more effective for complex point cloud data where local geometric structures are important.
@@
5class PointNet(nn.Module):
@@
8 self.conv1 = nn.Conv1d(3, 64, 1)
9 self.conv2 = nn.Conv1d(64, 128, 1)
10 self.conv3 = nn.Conv1d(128, 1024, 1)
11 self.fc1 = nn.Linear(1024, 512)
12 self.fc2 = nn.Linear(512, 256)
13 self.fc3 = nn.Linear(256, num_classes)
14 self.relu = nn.ReLU()
15 self.maxpool = nn.MaxPool1d(1024)
@@ 23.5 Transformer Models for Radar Data Processing
Transformers have become a dominant architecture in many fields of deep learning due to their ability to model long-range dependencies using the self-attention mechanism. Originally developed for natural language processing [51] (NLP), Transformer models are now being applied to radar data processing, where they excel at capturing complex relationships between data points. In this section, we explore how Transformer models can be used to process both radar-generated heatmaps and point clouds. By leveraging the self-attention mechanism, Transformers can improve target detection, classification, and segmentation tasks in radar systems.
@@ 23.5.1 Combining CNN and Transformer for Heatmaps
5class CNNTransformerModel(nn.Module):
@@
10 self.conv1 = nn.Conv2d(1, 32, kernel_size=3, padding=1) # input shape: (1, 100, 100)
11 self.pool1 = nn.MaxPool2d(2, 2) # output: (32, 50, 50)
@@
13 self.conv2 = nn.Conv2d(32, 64, kernel_size=3, padding=1) # output: (64, 50, 50)
14 self.pool2 = nn.MaxPool2d(2, 2) # output: (64, 25, 25)
@@
16 self.conv3 = nn.Conv2d(64, 128, kernel_size=3, padding=1) # output: (128, 25, 25)
@@
19 self.flatten = nn.Flatten()
@@
22 self.transformer = nn.TransformerEncoderLayer(
@@
27 self.fc = nn.Linear(128, num_classes)
@@ 23.5.2 PointNet++ and Transformer for Point Clouds
5class PointNetTransformer(nn.Module):
@@
9 self.conv1 = nn.Conv1d(3, 64, 1)
10 self.conv2 = nn.Conv1d(64, 128, 1)
11 self.conv3 = nn.Conv1d(128, 256, 1)
@@
14 self.multihead_attn = nn.MultiheadAttention(embed_dim=256, num_heads=8)
15 self.fc = nn.Linear(256, num_classes)
@@ 23.6 Conclusion
The combination of CNNs and Transformers for heatmap processing and PointNet++ with Transformers for point cloud processing provides a powerful approach for radar data analysis [69]. CNNs effectively extract local features, while Transformers capture global dependencies, enhancing the overall accuracy of radar data processing. Similarly, PointNet++ excels at local feature extraction in point clouds, and the integration of Transformers helps model complex relationships between distant points, improving the accuracy of classification, segmentation, and target detection tasks.
@@ 24.1 Importance of HAR in Radar Systems
Human Activity Recognition (HAR) has become an essential application in modern radar systems, particularly for domains such as security, health monitoring, and smart home environments. HAR systems use radar data to detect and classify human movements and activities, ranging from basic actions like walking or sitting to more complex patterns such as falls or gestures [52, 49]. Unlike traditional sensors like cameras, which are often limited by privacy concerns and environmental conditions, radar-based HAR offers unique advantages.
@@
The applications of radar-based HAR systems are vast. For example, in the field of security [31], HAR systems can monitor suspicious behaviors, alerting authorities to potential intruders. In healthcare [32], such systems can continuously track patients’ activities, identifying abnormal movements or falls. Smart homes [43] equipped with HAR can provide automated assistance for elderly or disabled individuals by recognizing activities such as walking, sitting, or standing up. These systems can enable automatic lighting, adjust temperature, or send alerts in emergency situations.
@@ 24.2.1 Data Noise and Signal Interference
Radar systems are inherently sensitive to noise [50], especially in environments with multiple moving objects or reflective surfaces. In a typical household or public setting, there are numerous potential sources of interference, including other moving objects, electronic devices, and reflective surfaces like walls or furniture. These factors can introduce noise and clutter into the radar signal, making it more difficult to accurately identify the human activity of interest.
@@ 24.2.4 High Dimensionality of Radar Data
Radar sensors generate high-dimensional data, often in the form of range-Doppler maps, micro-Doppler signatures [58], or point clouds. While these rich data representations provide valuable information about human activities, they also present challenges due to the high dimensionality. Processing and analyzing large amounts of radar data in real-time requires advanced techniques, such as deep learning, to extract meaningful features without overwhelming computational resources.
@@ 24.2.6 Example of Signal Noise in HAR
Let’s consider an example of how noise might affect radar-based HAR [28]. Suppose a radar system is tracking a person walking in a room, but there is a fan moving in the background. The fan’s motion generates a Doppler shift in the radar signal, which may interfere with the system’s ability to distinguish between the person walking and the fan’s movement.
@@ 25.1.1 Range-Doppler Maps and Spatial Features
The Range-Doppler Map is a 2D image [3] that contains rich information about the motion of objects relative to the radar. For human activity recognition, these maps capture movement patterns of body parts. CNNs can efficiently identify and learn these spatial features through convolutional layers, pooling layers, and activation functions.
@@
6class CNNModel(nn.Module):
@@
11 self.conv1 = nn.Conv2d(in_channels=1, out_channels=32, kernel_size=3, padding=1)
12 self.pool1 = nn.MaxPool2d(2, 2) # Pooling layer to reduce dimensions
@@
15 self.conv2 = nn.Conv2d(in_channels=32, out_channels=64, kernel_size=3, padding=1)
16 self.pool2 = nn.MaxPool2d(2, 2)
@@
19 self.conv3 = nn.Conv2d(in_channels=64, out_channels=128, kernel_size=3, padding=1)
20 self.pool3 = nn.MaxPool2d(2, 2)
@@
23 self.fc1 = nn.Linear(128 * 8 * 8, 128) # Assuming input image size is 64x64
@@
26 self.fc2 = nn.Linear(128, num_classes)
@@ 25.2 LSTM for Temporal Sequence Modeling
While CNNs are highly effective for extracting spatial features from individual radar images, they do not capture the temporal dynamics inherent in human activities. Many human actions (e.g., walking, sitting, standing) involve movement over time, and it is crucial to model these changes to recognize the activity correctly. Long Short-Term Memory [54] (LSTM) networks, a type of recurrent neural network (RNN), are designed to handle such temporal sequence data.
@@ 25.2.1 How LSTMs Handle Temporal Data
6class LSTMModel(nn.Module):
@@
11 self.lstm1 = nn.LSTM(input_size=input_size, hidden_size=hidden_size1, batch_first=True)
@@
14 self.lstm2 = nn.LSTM(input_size=hidden_size1, hidden_size=hidden_size2, batch_first=True)
@@
17 self.fc1 = nn.Linear(hidden_size2, 128)
@@
20 self.fc2 = nn.Linear(128, num_classes)
@@
45criterion = nn.CrossEntropyLoss() # For multi-class classification
@@ 25.3.1 Architecture of a CNN+LSTM Model
In this example, the radar images (RDMs) are first processed by the CNN, and the output features are reshaped and passed to the LSTM [71], which models the temporal dynamics. Finally, the output layer provides the activity classification.
@@ 25.4 Benefits of CNN+LSTM for HAR
Human Activity Recognition (HAR) is a complex task that involves detecting and identifying human movements using various data sources. In radar-based systems, point cloud data is an effective representation of human motion in 3D space [13]. The combination of PointNet for spatial feature extraction and LSTM (Long Short-Term Memory) networks for temporal sequence modeling provides a powerful solution for HAR using point cloud data.
@@ 26.1 PointNet for Point Cloud Feature Extraction
6class PointNet(nn.Module):
@@
9 self.mlp1 = nn.Sequential(
10 nn.Conv1d(3, 64, 1),
11 nn.BatchNorm1d(64),
12 nn.ReLU()
@@
14 self.mlp2 = nn.Sequential(
15 nn.Conv1d(64, 128, 1),
16 nn.BatchNorm1d(128),
17 nn.ReLU()
@@
19 self.mlp3 = nn.Sequential(
20 nn.Conv1d(128, 1024, 1),
21 nn.BatchNorm1d(1024),
22 nn.ReLU()
@@
24 self.fc1 = nn.Linear(1024, 512)
25 self.fc2 = nn.Linear(512, 256)
26 self.fc3 = nn.Linear(256, 128)
@@
$(batch\_size,3,num\_points)$. The PointNet extracts features from each point cloud using multi-layer perceptrons and max pooling.
@@ 26.2 LSTM for Temporal Dynamics in Point Clouds
3class LSTMModel(nn.Module):
@@
6 self.lstm = nn.LSTM(input_size, hidden_size, num_layers, batch_first=True)
7 self.fc = nn.Linear(hidden_size, num_classes)
@@ 26.3 Combining PointNet and LSTM for HAR
The combination of PointNet and LSTM allows for effective human activity recognition by simultaneously modeling spatial and temporal features. The PointNet model extracts spatial features from each point cloud frame, while the LSTM model captures the temporal dependencies between consecutive frames. This architecture is especially useful for radar-based HAR, where movements are captured as point cloud data over time [12].
@@
3class PointNetLSTM(nn.Module):
@@
8 self.fc = nn.Linear(128, num_classes)

### v2

@@ 23.2 Conclusion
Generating heatmaps from Range-Doppler Maps and using CNNs for target detection is an effective approach for radar data processing [de2020deep]. Heatmaps provide a visual representation of radar data, and CNNs automatically extract spatial features from the heatmaps, enabling the classification of targets in real-time radar applications. With the ability to handle noise and extract meaningful features, CNNs play a critical role in enhancing radar-based object detection systems.
@@ 23.3.1 Using CNN for Heatmap Classification
17class SimpleCNN($n$.Module):
@@
20 self.conv1 = $n$.Conv2d(1, 32, kernel_size=3, padding=1) # 1 input channel (grayscale), 32 output filters
21 self.pool = $n$.MaxPool2d(2, 2)
22 self.conv2 = $n$.Conv2d(32, 64, kernel_size=3, padding=1)
23 self.conv3 = $n$.Conv2d(64, 64, kernel_size=3, padding=1)
24 self.fc1 = $n$.Linear(64 * 25 * 25, 64) # Flatten after pooling
25 self.fc2 = $n$.Linear(64, 2) # Binary classification (target or noise)
@@
42criterion = $n$.CrossEntropyLoss() # For binary classification with logits
@@ 23.3.2 PointNet and PointNet++ for Point Cloud Processing
PointNet is a deep learning model that directly processes point cloud data without converting it into voxel grids or meshes [charles2017pointnet]. The architecture of PointNet is designed to handle unordered sets of points, making it ideal for processing radar-generated point clouds. The model extracts global features from the entire point cloud by applying a series of shared multi-layer perceptrons (MLPs) to each point, followed by a max pooling operation to aggregate the features.
@@
PointNet++ builds upon the success of PointNet by introducing a hierarchical structure that captures both global and local features [hao2023improved]. While PointNet extracts global features from the entire point cloud, PointNet++ divides the point cloud into smaller regions, applying PointNet to each region to extract local features. These local features are then combined to form a hierarchical representation of the point cloud, making PointNet++ more effective for complex point cloud data where local geometric structures are important.
@@
5class PointNet($n$.Module):
@@
8 self.conv1 = $n$.Conv1d(3, 64, 1)
9 self.conv2 = $n$.Conv1d(64, 128, 1)
10 self.conv3 = $n$.Conv1d(128, 1024, 1)
11 self.fc1 = $n$.Linear(1024, 512)
12 self.fc2 = $n$.Linear(512, 256)
13 self.fc3 = $n$.Linear(256, num_classes)
14 self.relu = $n$.ReLU()
15 self.maxpool = $n$.MaxPool1d(1024)
@@ 23.5 Transformer Models for Radar Data Processing
Transformers have become a dominant architecture in many fields of deep learning due to their ability to model long-range dependencies using the self-attention mechanism. Originally developed for natural language processing [sengupta2022mmpose] (NLP), Transformer models are now being applied to radar data processing, where they excel at capturing complex relationships between data points. In this section, we explore how Transformer models can be used to process both radar-generated heatmaps and point clouds. By leveraging the self-attention mechanism, Transformers can improve target detection, classification, and segmentation tasks in radar systems.
@@ 23.5.1 Combining CNN and Transformer for Heatmaps
5class CNNTransformerModel($n$.Module):
@@
10 self.conv1 = $n$.Conv2d(1, 32, kernel_size=3, padding=1) # input shape: (1, 100, 100)
11 self.pool1 = $n$.MaxPool2d(2, 2) # output: (32, 50, 50)
@@
13 self.conv2 = $n$.Conv2d(32, 64, kernel_size=3, padding=1) # output: (64, 50, 50)
14 self.pool2 = $n$.MaxPool2d(2, 2) # output: (64, 25, 25)
@@
16 self.conv3 = $n$.Conv2d(64, 128, kernel_size=3, padding=1) # output: (128, 25, 25)
@@
19 self.flatten = $n$.Flatten()
@@
22 self.transformer = $n$.TransformerEncoderLayer(
@@
27 self.fc = $n$.Linear(128, num_classes)
@@ 23.5.2 PointNet++ and Transformer for Point Clouds
5class PointNetTransformer($n$.Module):
@@
9 self.conv1 = $n$.Conv1d(3, 64, 1)
10 self.conv2 = $n$.Conv1d(64, 128, 1)
11 self.conv3 = $n$.Conv1d(128, 256, 1)
@@
14 self.multihead_attn = $n$.MultiheadAttention(embed_dim=256, num_heads=8)
15 self.fc = $n$.Linear(256, num_classes)
@@ 23.6 Conclusion
The combination of CNNs and Transformers for heatmap processing and PointNet++ with Transformers for point cloud processing provides a powerful approach for radar data analysis [zhou2022centerformer]. CNNs effectively extract local features, while Transformers capture global dependencies, enhancing the overall accuracy of radar data processing. Similarly, PointNet++ excels at local feature extraction in point clouds, and the integration of Transformers helps model complex relationships between distant points, improving the accuracy of classification, segmentation, and target detection tasks.
@@ 24.1 Importance of HAR in Radar Systems
Human Activity Recognition (HAR) has become an essential application in modern radar systems, particularly for domains such as security, health monitoring, and smart home environments. HAR systems use radar data to detect and classify human movements and activities, ranging from basic actions like walking or sitting to more complex patterns such as falls or gestures [ShahzSec13, saebaBio12]. Unlike traditional sensors like cameras, which are often limited by privacy concerns and environmental conditions, radar-based HAR offers unique advantages.
@@
The applications of radar-based HAR systems are vast. For example, in the field of security [lee2015radar], HAR systems can monitor suspicious behaviors, alerting authorities to potential intruders. In healthcare [li2010recent], such systems can continuously track patients’ activities, identifying abnormal movements or falls. Smart homes [perez2022detecting] equipped with HAR can provide automated assistance for elderly or disabled individuals by recognizing activities such as walking, sitting, or standing up. These systems can enable automatic lighting, adjust temperature, or send alerts in emergency situations.
@@ 24.2.1 Data Noise and Signal Interference
Radar systems are inherently sensitive to noise [schmitt2009radar], especially in environments with multiple moving objects or reflective surfaces. In a typical household or public setting, there are numerous potential sources of interference, including other moving objects, electronic devices, and reflective surfaces like walls or furniture. These factors can introduce noise and clutter into the radar signal, making it more difficult to accurately identify the human activity of interest.
@@ 24.2.4 High Dimensionality of Radar Data
Radar sensors generate high-dimensional data, often in the form of range-Doppler maps, micro-Doppler signatures [tahmoush2015review], or point clouds. While these rich data representations provide valuable information about human activities, they also present challenges due to the high dimensionality. Processing and analyzing large amounts of radar data in real-time requires advanced techniques, such as deep learning, to extract meaningful features without overwhelming computational resources.
@@ 24.2.6 Example of Signal Noise in HAR
Let’s consider an example of how noise might affect radar-based HAR [kim2022radar]. Suppose a radar system is tracking a person walking in a room, but there is a fan moving in the background. The fan’s motion generates a Doppler shift in the radar signal, which may interfere with the system’s ability to distinguish between the person walking and the fan’s movement.
@@ 25.1.1 Range-Doppler Maps and Spatial Features
The Range-Doppler Map is a 2D image [altmann2018deep] that contains rich information about the motion of objects relative to the radar. For human activity recognition, these maps capture movement patterns of body parts. CNNs can efficiently identify and learn these spatial features through convolutional layers, pooling layers, and activation functions.
@@
6class CNNModel($n$.Module):
@@
11 self.conv1 = $n$.Conv2d(in_channels=1, out_channels=32, kernel_size=3, padding=1)
12 self.pool1 = $n$.MaxPool2d(2, 2) # Pooling layer to reduce dimensions
@@
15 self.conv2 = $n$.Conv2d(in_channels=32, out_channels=64, kernel_size=3, padding=1)
16 self.pool2 = $n$.MaxPool2d(2, 2)
@@
19 self.conv3 = $n$.Conv2d(in_channels=64, out_channels=128, kernel_size=3, padding=1)
20 self.pool3 = $n$.MaxPool2d(2, 2)
@@
23 self.fc1 = $n$.Linear(128 * 8 * 8, 128) # Assuming input image size is 64x64
@@
26 self.fc2 = $n$.Linear(128, num_classes)
@@ 25.2 LSTM for Temporal Sequence Modeling
While CNNs are highly effective for extracting spatial features from individual radar images, they do not capture the temporal dynamics inherent in human activities. Many human actions (e.g., walking, sitting, standing) involve movement over time, and it is crucial to model these changes to recognize the activity correctly. Long Short-Term Memory [shrestha2020continuous] (LSTM) networks, a type of recurrent neural network (RNN), are designed to handle such temporal sequence data.
@@ 25.2.1 How LSTMs Handle Temporal Data
6class LSTMModel($n$.Module):
@@
11 self.lstm1 = $n$.LSTM(input_size=input_size, hidden_size=hidden_size1, batch_first=True)
@@
14 self.lstm2 = $n$.LSTM(input_size=hidden_size1, hidden_size=hidden_size2, batch_first=True)
@@
17 self.fc1 = $n$.Linear(hidden_size2, 128)
@@
20 self.fc2 = $n$.Linear(128, num_classes)
@@
45criterion = $n$.CrossEntropyLoss() # For multi-class classification
@@ 25.3.1 Architecture of a CNN+LSTM Model
In this example, the radar images (RDMs) are first processed by the CNN, and the output features are reshaped and passed to the LSTM [zhu2020hybrid], which models the temporal dynamics. Finally, the output layer provides the activity classification.
@@ 25.4 Benefits of CNN+LSTM for HAR
Human Activity Recognition (HAR) is a complex task that involves detecting and identifying human movements using various data sources. In radar-based systems, point cloud data is an effective representation of human motion in 3D space [ding2021radar]. The combination of PointNet for spatial feature extraction and LSTM (Long Short-Term Memory) networks for temporal sequence modeling provides a powerful solution for HAR using point cloud data.
@@ 26.1 PointNet for Point Cloud Feature Extraction
6class PointNet($n$.Module):
@@
9 self.mlp1 = $n$.Sequential(
10 $n$.Conv1d(3, 64, 1),
11 $n$.BatchNorm1d(64),
12 $n$.ReLU()
@@
14 self.mlp2 = $n$.Sequential(
15 $n$.Conv1d(64, 128, 1),
16 $n$.BatchNorm1d(128),
17 $n$.ReLU()
@@
19 self.mlp3 = $n$.Sequential(
20 $n$.Conv1d(128, 1024, 1),
21 $n$.BatchNorm1d(1024),
22 $n$.ReLU()
@@
24 self.fc1 = $n$.Linear(1024, 512)
25 self.fc2 = $n$.Linear(512, 256)
26 self.fc3 = $n$.Linear(256, 128)
@@
(batch_size,3,num_points)(batch\_size,3,num\_points). The PointNet extracts features from each point cloud using multi-layer perceptrons and max pooling.
@@ 26.2 LSTM for Temporal Dynamics in Point Clouds
3class LSTMModel($n$.Module):
@@
6 self.lstm = $n$.LSTM(input_size, hidden_size, num_layers, batch_first=True)
7 self.fc = $n$.Linear(hidden_size, num_classes)
@@ 26.3 Combining PointNet and LSTM for HAR
The combination of PointNet and LSTM allows for effective human activity recognition by simultaneously modeling spatial and temporal features. The PointNet model extracts spatial features from each point cloud frame, while the LSTM model captures the temporal dependencies between consecutive frames. This architecture is especially useful for radar-based HAR, where movements are captured as point cloud data over time [ding2023sparsity].
@@
3class PointNetLSTM($n$.Module):
@@
8 self.fc = $n$.Linear(128, num_classes)